# HTD-Mamba: Efficient Hyperspectral Target Detection with Pyramid State Space Model

### v1

@@ II-E Spectral Contrastive Head
To maximize the similarity of positive samples while minimizing the similarity of negative samples, contrastive learning is required. In this paper, the feature vector pairs $({\bf x^{\prime\prime}},{\bf y^{\prime\prime}})$ from the same location construct the positive samples, and those from different locations construct the negative ones. For example, given a batch of size $P$ and a specific sample ${\bf x^{\prime\prime}}^{(i)}$, there are $P$ pairs in total, where $({\bf x^{\prime\prime}}^{(i)},{\bf y^{\prime\prime}}^{(i)})$ is the only positive pair while the others are negative pairs. Thus, the loss for ${\bf x^{\prime\prime}}^{(i)}$ can be expressed as
$l({\bf x^{\prime\prime}}^{(i)},{\bf y^{\prime\prime}}^{(i)})=-{\rm log}\frac{{\rm exp}(s({\bf x^{\prime\prime}}^{(i)},{\bf y^{\prime\prime}}^{(i)})/\alpha)}{\sum_{j=1}^{P}{{\rm exp}(s({\bf x^{\prime\prime}}^{(i)},{\bf y^{\prime\prime}}^{(j)})/\alpha)}}$ (19)
where $s({\bf a},{\bf b})=({\bf a}^{T}{\bf b})/(\|{\bf a}\|_{2}\cdot\|{\bf b}\|_{2})$ denotes the pairwise similarity, and $\alpha$ is the temperature parameter to control the degree of attention to negative pairs. Taking into account learning the similarity of all the positive pairs within a batch, the spectral contrastive loss can be expressed as
$L_{\rm contrast}=\frac{1}{P}\sum_{i=1}^{P}l({\bf x^{\prime\prime}}^{(i)},{\bf y^{\prime\prime}}^{(i)}).$ (20)
@@ III-A1 Datasets
III-A2 Quality Metrics
To assess the quality of the detection map, both qualitative and quantitative metrics are adopted. Specifically, the visualized detection map, receiver operating characteristic (ROC) curve, and background-target separability diagram are utilized for qualitative assessment. The ROC curve is a function of false alarm rate $P_{f}$ and detection probability $P_{d}$ under different thresholds $\tau$. For comprehensive analysis, 3D ROC curves [23] are adopted, where three unfolded 2D ROC curves of $(P_{f},P_{d})$, $(P_{f},\tau)$, and $(P_{d},\tau)$ are obtained to measure detection effectiveness, target detectability, and background suppression, respectively. The background-target separability map is plotted using the Box–Whisker function to show the distinction between background and target. For quantitative evaluation, three areas under the ROC curves (AUCs) are applied, i.e., ${\rm AUC}(P_{f},P_{d})$, ${\rm AUC}(\tau,P_{d})$, and ${\rm AUC}(\tau,P_{f})$. In addition, two composite metrics are also calculated to evaluate overall quality:
${\rm AUC_{SNPR}}={\rm AUC}(\tau,P_{d})/{\rm AUC}(\tau,P_{f}).$ (22)
The detailed properties of these five metrics are listed in Table I.
@@ III-A3 Implementation Details
The proposed HTD-Mamba is implemented using Python 3.10.4 and PyTorch 2.0.1, with an Intel Core i9-10900X CPU, 32 GB of RAM, and an NVIDIA TITAN RTX with 24 GB of GPU memory. The batch size $P$ is set to 80. For each batch, the patch size surrounding the center pixel is set to $11\times 1$ 11, $13\times 13$, $11\times 1$ 11, and $5\times 5$ for San Diego I, San Diego II, Los Angeles, and Pavia, respectively. In the group-wise spectral embedding, the group length $m$ is set to 30, 5, 5, and 15 for the four datasets, respectively. The embedding dimension $N$ is set to 16 for all four datasets. In the feature extraction and projection blocks, the depth of the pyramid SSM is set to 1, the hidden state size $D$ is set to 16, and the dimension of the final feature vectors $d$ is set to 32 for all four datasets. For the contrastive loss, the temperature parameter $\alpha$ is fixed at 0.1 for all four datasets. The AdamW optimizer is selected for training, where the weight-decay parameter is set to $10^{-4}$. The initial learning rate is set to $10^{-4}$, which increases via a linear warmup scheduler in the first $10\%$ epochs and then decreases via a cosine scheduler. The total number of epochs is 200. In the background suppression process, the parameter $\delta$ is set to 0.1 for all four datasets.
@@ III-B4 Network Depth $l$
The network depth $l$ represents the number of deep layers in the pyramid SSM. To investigate its impact on detection performance, we compared the AUC scores for network depths ranging from 1 to 5. The experimental results are shown in Fig. 4(d). As shown in the figure, detection accuracy remains stable or slightly decreases with increasing $l$. This is because the pyramid SSM inherently possesses the ability to extract global spectral features with multi-resolution using shallow layers. When the number of layers increases, the model may overfit, leading to decreased accuracy and increased computational complexity. Therefore, we set $l$ to 1 for all datasets to ensure a lightweight model while maintaining detection effectiveness.
@@ III-C Performance Comparison
For fair comparison, the key parameters of the competing methods are tuned to optimal levels. Specifically, for the OSP detector, the automatic target generation process (ATGP) [31] is applied to obtain the background subspace, where the number of bases is selected as 22, 2, 1, and 1 for San Diego I, San Diego II, Los Angeles, and Pavia, respectively. In CSCR, the dual window size $(w_{out},w_{in})$ is set to $(13,11)$, $(15,3)$, $(15,13)$, and $(15,13)$ for the four datasets, respectively. Additionally, the two regularization parameters $(\lambda,\beta)$ are set to $(0.1,1)$ for all datasets. For CTTD, the vertical division for spatial disassembly $ver$ and the horizontal division for spectral disassembly $hor$, are set to $(18,2)$, $(14,4)$, $(3,4)$, and $(18,7)$ for the four datasets, respectively. For BLTSC, the binarization threshold $k$ in coarse detection is fixed at 0.15, the learning rate is $10^{-4}$, the batch size is 64, the number of epochs is 500, and the parameter $\lambda$ in nonlinear background suppression is set to $10^{3}$, $10^{4}$, $10^{4}$, and 10 for the four datasets, respectively. For MLSN, the batch size, learning rate, and epoch are set to 128, $10^{-3}$, and 50, respectively. In the process of guided image filtering, the local window radius is set to 2, and the penalty value is set to 0.04 for all datasets. For OS-VAE, the number of hidden nodes, the network depth, the regularization parameter $\rho$, and the background suppression parameter $\alpha$ are respectively set to 30, 2, $10^{-5}$, and 0.1 for all datasets. For TSTTD, the mixture ratio of background in synthesizing target samples is randomly set between 0 and 0.1. During the training process, the batch size is set to 64, the learning rate is $10^{-4}$, and the total number of epochs is 20.
@@
To further compare the background-target separation performance of the competing methods, Fig. 6 shows the separation diagrams on the four datasets. It is evident that our proposed HTD-Mamba achieves the most satisfactory performance, as there is significant distance and minimal overlap between the target and background boxes across all four datasets. Although BLTSC, TSTTD, and HTD-Mamba all achieve superior background suppression performance with small-sized background boxes, our proposed HTD-Mamba demonstrates a more promising target-background separation. Overall, HTD-Mamba consistently produces satisfactory visual detection maps, achieves outstanding quantitative results, and demonstrates promising target-background separation capabilities across all four datasets.
@@ III-D1 Effectiveness of SESA
The proposed SESA technique encodes surrounding contextual information to create a new spectral view of the central pixel, providing sufficient view pairs for contrastive learning. To verify its effectiveness, we compare SESA with the no-transform strategy (referred to as self-similarity) and several spectral data augmentation methods, including Dropout[32], Gaussian noise, variational autoencoder (VAE)[33], and adversarial autoencoder (AAE)[34]. Specifically, self-similarity treats each pixel and its copy as a positive pair, while different pixels are treated as negative pairs. Dropout randomly deactivates a certain percentage (0.5 in this paper) of the bands of each pixel to create a new spectral view, which is then paired with the original pixel to construct the positive sample. The Gaussian noise method adds random noise (20 dB in this paper) to the pixel to generate a degraded view simulating complex scenes. Both VAE and AAE are autoregressive methods that make the latent space approach prior Gaussian distributions by maximizing the variational lower bound or through adversarial training. Pixels can be paired with their regressed variants generated by VAE or AAE to construct view pairs for training. We adopt the same autoencoder architecture as used in [15] for both VAE and AAE, which includes a spectral residual channel attention module and a simple discriminator composed of two FC layers for AAE. For a fair comparison, the detected pixel is directly paired with the target spectrum during inference.
@@ III-D2 Effectiveness of Pyramid SSM
To further verify the superiority of the pyramid SSM, we compare it with several mainstream backbones, including MLP [35], long short-term memory (LSTM) [36], ResNet [37], vision Transformer (ViT) [38], Mamba [21], feature pyramid network (FPN) [39], U-Net [40], and pyramid vision Transformer (PVT) [41]. Among these methods, MLP, LSTM, ResNet, ViT, and Mamba are single-scale based models, while FPN, UNet, and PVT are multi-scale based models considering different resolutions. Since ResNet, ViT, FPN, UNet, and PVT are originally based on RGB images, we modified them into the 1D form for spectral processing. For a fair comparison, the parameter capacity of all the competing models is set to roughly the same by adjusting the network depth. The comparison results are shown in Table V. It can be found that all the competing models achieve promising performance, but our pyramid SSM obtains the best results on all four datasets. Additionally, with similar parameters, pyramid SSM requires relatively low floating point operations (FLOPs), and the running time is also within a feasible range for practical applications. Therefore, the proposed pyramid SSM is effective and efficient for extracting discriminative intrinsic features in HTD due to its consideration of multiresolution global correlations and the hardware parallel scanning mechanism.

### v2

@@ II-E Spectral Contrastive Head
To maximize the similarity of positive samples while minimizing the similarity of negative samples, contrastive learning is required. In this paper, the feature vector pairs $({\bf x^{\prime\prime}},{\bf y^{\prime\prime}})$ from the same location construct the positive samples, and those from different locations construct the negative ones. For example, given a batch of $({\bf x^{\prime\prime}},{\bf y^{\prime\prime}})$, for a specific sample ${\bf x^{\prime\prime}}^{(i)}$, there are $P$ pairs in total, where $({\bf x^{\prime\prime}}^{(i)},{\bf y^{\prime\prime}}^{(i)})$ is the only positive pair while the others are negative pairs. Thus, the loss for ${\bf x^{\prime\prime}}^{(i)}$ can be expressed as
$\ell({\bf x^{\prime\prime}}^{(i)},{\bf y^{\prime\prime}}^{(i)})=-{\rm log}\frac{{\rm exp}(c({\bf x^{\prime\prime}}^{(i)},{\bf y^{\prime\prime}}^{(i)})/\alpha)}{\sum_{j=1}^{P}{{\rm exp}(c({\bf x^{\prime\prime}}^{(i)},{\bf y^{\prime\prime}}^{(j)})/\alpha)}}$ (18)
where $\alpha$ is the temperature parameter to control the degree of attention to negative pairs. Taking into account learning the similarity of all the positive pairs within a batch, the spectral contrastive loss can be expressed as
$L_{\rm contrast}=\frac{1}{P}\sum_{i=1}^{P}\ell({\bf x^{\prime\prime}}^{(i)},{\bf y^{\prime\prime}}^{(i)}).$ (19)
@@ III-A1 Datasets
III-A2 Quality Metrics
To assess the quality of the detection map, both qualitative and quantitative metrics are adopted. Specifically, the visualized detection map, receiver operating characteristic (ROC) curve, and background-target separability diagram are utilized for qualitative assessment. The ROC curve is a function of false alarm rate $P_{f}$ and detection probability $P_{d}$ under different thresholds $\tau$. For comprehensive analysis, 3D ROC curves [37] are adopted, where three unfolded 2D ROC curves of $(P_{f},P_{d})$, $(P_{f},\tau)$, and $(P_{d},\tau)$ are obtained to measure detection effectiveness, target detectability, and background suppression, respectively. The background-target separability diagram is plotted using the Box–Whisker function to show the distinction between background and target. For quantitative evaluation, three areas under the ROC curves (AUCs) are applied, i.e., ${\rm AUC}(P_{f},P_{d})$, ${\rm AUC}(\tau,P_{d})$, and ${\rm AUC}(\tau,P_{f})$. In addition, two composite metrics are also calculated to evaluate overall quality:
${\rm AUC_{SNPR}}={\rm AUC}(\tau,P_{d})/{\rm AUC}(\tau,P_{f}).$ (21)
The detailed properties of these five metrics are listed in Table I.
@@ III-A3 Implementation Details
The proposed HTD-Mamba is implemented using Python 3.10.4 and PyTorch 2.0.1, with an Intel Core i9-10900X CPU, 32 GB of RAM, and an NVIDIA TITAN RTX with 24 GB of GPU memory. The batch size $P$ is set to 80. For each batch, the patch size surrounding the center pixel is set to $11\times 1$ 11, $13\times 13$, $11\times 1$ 11, and $5\times 5$ for San Diego I, San Diego II, Los Angeles, and Pavia, respectively. In the group-wise spectral embedding, the group length $m$ is set to 30, 5, 5, and 15 for the four datasets, respectively. The embedding size $N$ is set to 16 for all four datasets. In the feature extraction and projection blocks, the depth of the pyramid SSM is set to 1, the hidden state size $D$ is set to 16, and the dimension of the final feature vectors $d$ is set to 32 for all four datasets. For the contrastive loss, the temperature parameter $\alpha$ is fixed at 0.1 for all four datasets. The AdamW optimizer is selected for training, where the weight-decay parameter is set to $10^{-4}$. The initial learning rate is set to $10^{-4}$, which increases via a linear warmup scheduler in the first $10\%$ epochs and then decreases via a cosine scheduler. The total number of epochs is 200. In the background suppression process, the parameter $\delta$ is set to 0.1 for all four datasets.
@@ III-B4 Network Depth $l$
The network depth $l$ represents the number of deep layers in the pyramid SSM. To investigate its impact on detection performance, we compared the AUC scores for network depths ranging from 1 to 5. The experimental results are shown in Fig. 4(d). As shown in the figure, detection accuracy remains stable or slightly decreases with increasing $l$. This is because the pyramid SSM inherently possesses the ability to extract global spectral features with multiresolution using shallow layers. When the number of layers increases, the model may overfit, leading to decreased accuracy and increased computational complexity. Therefore, we set $l$ to 1 for all datasets to ensure a lightweight model while maintaining detection effectiveness.
@@ III-C Performance Comparison
For a fair comparison, the key parameters of the competing methods are tuned to optimal levels. Specifically, for the OSP detector, the automatic target generation process (ATGP) [40] is applied to obtain the background subspace, where the number of bases is selected as 22, 2, 1, and 1 for San Diego I, San Diego II, Los Angeles, and Pavia, respectively. In CSCR, the dual window size $(w_{out},w_{in})$ is set to $(13,11)$, $(15,3)$, $(15,13)$, and $(15,13)$ for the four datasets, respectively. Additionally, the two regularization parameters $(\lambda,\beta)$ are set to $(0.1,1)$ for all datasets. For CTTD, the vertical division for spatial disassembly $ver$ and the horizontal division for spectral disassembly $hor$, are set to $(18,2)$, $(14,4)$, $(3,4)$, and $(18,7)$ for the four datasets, respectively. For BLTSC, the binarization threshold $k$ in coarse detection is fixed at 0.15, the learning rate is $10^{-4}$, the batch size is 64, the number of epochs is 500, and the parameter $\lambda$ in nonlinear background suppression is set to $10^{3}$, $10^{4}$, $10^{4}$, and 10 for the four datasets, respectively. For MLSN, the batch size, learning rate, and epoch are set to 128, $10^{-3}$, and 50, respectively. In the process of guided image filtering, the local window radius is set to 2, and the penalty value is set to 0.04 for all datasets. For OS-VAE, the number of hidden nodes, the network depth, the regularization parameter $\rho$, and the background suppression parameter $\alpha$ are respectively set to 30, 2, $10^{-5}$, and 0.1 for all datasets. For TSTTD, the mixture ratio of background in synthesizing target samples is randomly set between 0 and 0.1. During the training process, the batch size is set to 64, the learning rate is $10^{-4}$, and the total number of epochs is 20.
@@
To further compare the background-target separation performance of the competing methods, Fig. 6 shows the separation diagrams for the four datasets. It is evident that our proposed HTD-Mamba achieves the most satisfactory performance, as there is significant distance and minimal overlap between the target and background boxes across all four datasets. Although BLTSC, TSTTD, and HTD-Mamba all achieve superior background suppression performance with small-sized background boxes, our proposed HTD-Mamba demonstrates a more promising target-background separation. Overall, HTD-Mamba consistently produces satisfactory visual detection maps, achieves outstanding quantitative results, and demonstrates promising target-background separation capabilities across all four datasets.
@@ III-D1 Effectiveness of SESA
The proposed SESA technique encodes surrounding contextual information to create a new spectral view of the center pixel, providing sufficient view pairs for contrastive learning. To verify its effectiveness, we compare SESA with the no-transform strategy (referred to as self-similarity) and several spectral data augmentation methods, including Dropout[41], Gaussian noise, variational autoencoder (VAE)[42], and adversarial autoencoder (AAE)[43]. Specifically, self-similarity treats each pixel and its copy as a positive pair, while different pixels are treated as negative pairs. Dropout randomly deactivates a certain percentage (0.5 in this paper) of the bands of each pixel to create a new spectral view, which is then paired with the original pixel to construct the positive sample. The Gaussian noise method adds random noise (20 dB in this paper) to the pixel to generate a degraded view simulating complex scenes. Both VAE and AAE are autoregressive methods that encode the spectral information into a latent space, either by maximizing the variational lower bound or through adversarial training. Pixels can be paired with their regressed variants generated by VAE or AAE to construct view pairs for training. We adopt the same autoencoder architecture used in [27] for both VAE and AAE, which includes a spectral residual channel attention module and a simple discriminator composed of two FC layers for AAE. For a fair comparison, the detected pixel is directly paired with the target spectrum during inference.
@@ III-D2 Effectiveness of Pyramid SSM
To further verify the superiority of the pyramid SSM, we compare it with several mainstream backbones, including MLP [44], long short-term memory (LSTM) [45], residual network (ResNet) [46], vision Transformer (ViT) [47], Mamba [33], feature pyramid network (FPN) [48], U-Net [49], and pyramid vision Transformer (PVT) [50]. Among these methods, MLP, LSTM, ResNet, ViT, and Mamba are single-scale based models, while FPN, UNet, and PVT are multi-scale based models considering different resolutions. Since ResNet, ViT, FPN, UNet, and PVT are originally based on RGB images, we modified them into the 1D form for spectral processing. For a fair comparison, the parameter capacity of all the competing models is set to roughly the same by adjusting the network depth. The comparison results are shown in Table V. It can be found that all the competing models achieve promising performance, but our pyramid SSM obtains the best results on all four datasets. Additionally, with similar parameters, pyramid SSM requires relatively low floating point operations (FLOPs), and the running time is also within a feasible range for practical applications. Therefore, the proposed pyramid SSM is effective and efficient for extracting discriminative intrinsic features in HTD due to its consideration of multiresolution global correlations and the hardware parallel scanning mechanism.